OpenStudy (anonymous):

Write the equation of the line that passes through the points (–1, –4) and (2, 5). Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did. PLEEEAASEEE HELP ME ON THIS IT IS SOOOOOOO CONFUSINGGG :'(

hartnn (hartnn):

First find the slope of line: The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

5-(-4)/2-(-1) 9/3 3 is the slope now you can plug other values in to the point slope form y-y1=m(x-x1)

OpenStudy (anonymous):

y-5=3(x-2) y-5=3x-6 y=3x-1 y-(-4)=3(x-(-1)) so y+4=3x+3 y=3x-1 so ultimately no, the point wouldn't matter because the equations are the same

hero (hero):

Given: (–1, –4) and (2, 5) The question asks you to write the equation of a line. The formula for equation of a line is y = mx + b. Step 1: Insert each point one at a time into the formula y = mx + b to get: 5 = 2m + b -4 = -1m + b Step 2: Isolate b in each equation b = 5 - 2m b = m - 4 set b = b 5 - 2m = m - 4 Place like terms on the same side 5 + 4 = 2m + m 9 = 3m 9/3 = m 3 = m Find b: b = m - 4 b = 3 - 4 b = -1 Write the equation of a line y = mx + b by including the values found for m and b y = 3x - 1 You do not need either point in the final answer